OpenStudy (anonymous):

A small combination lock on a suitcase has 6 wheels, each labeled with the 10 digits 0 to 9. How many digit combinations are possible if no digit is repeated? If digits can be repeated? If successive digits must be different?

OpenStudy (anonymous):

if no digit is repeated = 10!, If digits can be repeated =10^9

OpenStudy (valpey):

With six wheels question 1 would be 10!/(10-6)! = 10!/4! = 10*9*8*7*6*5 Question 2 would be 10^6 = One million Question 3 would be 10*9^5